OpenStudy (anonymous):

Solve for x, a) log x= 3 b) log base 3 (x+2) = 1.732 c) log base 2 (x/10) = -0.671

OpenStudy (anonymous):

\[\log(x)\iff x=10^3\]

OpenStudy (anonymous):

\[\log_3(x+2)=1.732\iff x+2= 3^{1.732}\iff x=3^{1.732}-2\]

OpenStudy (anonymous):

ah thank you!!

OpenStudy (anonymous):

Exponentiate each side. a) log x=3 is equivalent to \[10^{3}=x=1000\]. Remember, a logarithm is an exponent! b) \[\log _{3}(x+2)=1.732\] written in exponential form, is \[3^{1.732}=x+2\] Therefore, \[x=3^{1.732}-2\]. Same for part c. Hope this helps

OpenStudy (anonymous):

ok.. so c would be... hmm x/10= 3^-0.671 x= 3^-0.671/ 10 ??? or is that completely wrong? haha

OpenStudy (anonymous):

i mean.. x= 3^-0.671*10

OpenStudy (anonymous):

damn it am stuck haha